bralyn:

The square root of 25 is rational. False True

Tonycoolkid21:

Is Square Root of 25 Rational or Irrational? A rational number can be expressed in the form of p/q. Because √25 = 5 and 5 can be written in the form of a fraction 5/1. It proves that √25 is rational.

bizaardvark:

This should help: A rational number can be expressed in the form of p/q. Because √25 = 5 and 5 can be written in the form of a fraction 5/1. It proves that √25 is rational. https://www.cuemath.com/algebra/square-root-of-25/#:~:text=25%20is%205.-,Is%20Square%20Root%20of%2025%20Rational%20or%20Irrational%3F,that%20%E2%88%9A25%20is%20rational.

bralyn:

@bizaardvark wrote:
This should help: A rational number can be expressed in the form of p/q. Because √25 = 5 and 5 can be written in the form of a fraction 5/1. It proves that √25 is rational. https://www.cuemath.com/algebra/square-root-of-25/#:~:text=25%20is%205.-,Is%20Square%20Root%20of%2025%20Rational%20or%20Irrational%3F,that%20%E2%88%9A25%20is%20rational.
didint help

bralyn:

do u know the anser

bralyn:

just asking u do not have to give it to me

bizaardvark:

Yes it tells you the answer in what i put up there

itsmehjay:

Do you know what rational and irrational muners are?

itsmehjay:

@itsmehjay wrote:
Do you know what rational and irrational muners are?
numbers* e.e